Direct Materials Variances Bellingham Company produces a product that requires 11 standard pounds per unit. The...

  1. Direct Materials Variances

    Bellingham Company produces a product that requires 11 standard pounds per unit. The standard price is $7.5 per pound. If 3,300 units used 37,800 pounds, which were purchased at $7.12 per pound, what is the direct materials (a) price variance, (b) quantity variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    a. Direct materials price variance $
    b. Direct materials quantity variance $
    c. Direct materials cost variance $
Answer #1

Materials Price Variance = (SP-AP) *AQ

= (7.5 - 7.12) *37800

= -14364 Favourable

Materials Quantity Variance = (SQ-AQ) *SP

= (3300*11 - 37800)*7.5

= 11,250 Unfavorable

Materials cost variance = 14364 F + 11250 U

= -3114 Favourable